# A modern perspective on rational homotopy theory

### v1

@@ 9.2. The construction of local spheres
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. To salvage our function complex characterization of $M$-local spaces in the absence of a universal property for $M$-localizatrion, we need the details of the CW complex construction of the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ from [FHT01, Chapter 9(a)].
Let $\mathbb{N}$ denote the set of positive integers. We enumerate
$M\cap\mathbb{N}=\left\{k_{1},k_{2},k_{3},\dots\right\}.$
We denote the $(n+1)$-disk by $D^{n+1}$. The $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ is constructed as the pushout
@@
[…]
The map
$(-k_{j})\colon S^{n}\rightarrow(S^{n})_{j}$
is a self-map of the $n$-sphere $S^{n}$ that has degree $(-k_{j})$. In other words, the $1$-connected CW complex
[…]
is the mapping telescope of the sequence of maps
$S^{n}\xrightarrow{k_{1}}S^{n}\xrightarrow{k_{2}}S^{n}\xrightarrow{k_{3}}\cdots.$
9.3. A characterization of local spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We use the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ from Section 9.2, including our notation, to prove the following key lemma.
Lemma 9.5.
@@ 9.3. A characterization of local spaces
Given a map $f\colon S^{n}\rightarrow Y$, we know that $[f]$ is an element of the $M^{-1}\mathbb{Z}$-module $\pi_{n}(Y)$. Thus, there exists a unique element $[u]$ of $\pi_{n}(Y)$ such that
$[f]=k_{1}\cdot[u],$
and the representative $u\colon S^{n}\rightarrow Y$ of $[u]$ is unique up to homotopy. Using the notation in our construction of
[…]
we extend $f$ from $S^{n}=(S^{n})_{0}$ to $(S^{n})_{0}\vee(S^{n})_{1}$ uniquely up to homotopy by setting
$g|_{(S^{n})_{1}}=u.$
As
$[f]-\left(k_{1}\cdot[u]\right)=0\in\pi_{n}(Y),$
we can further extend $f$ to
@@
In particular, we need the following consequence of Lemma 9.5 to give a characterization of $M$-local spaces.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. Let $Y$ be an $M$-local space. Then, every map $f\colon S^{n}\rightarrow Y$ extends uniquely up to homotopy to a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$ such that the diagram
@@
commutes.
We salvage our characterization of $M$-local spaces for the left Bousfield localization in the next section.
Proposition 9.7.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
Proof.
If $Y$ is $M$-local, then Corollary 9.6 implies that, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
is a weak homotopy equivalence. The converse implication is established as in the proof of Proposition 6.8. ∎
10. The family of model structures on simplicial sets
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize the model category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets to obtain a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ whose fibrant objects are the $M$-local ones. In particular, we produce a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ whose fibrant objects are the rational simplicial sets. Thus, we partially extend Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets. Our family of model structures is meant to axiomatize the generalization of rational homotopy theory that we studied in Section 9.
10.1. The set for the left Bousfield localization
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
$T_{M^{-1}\mathbb{Z}}=\left\{E_{1}\operatorname{Sing}(S^{n})\rightarrow E_{1}\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}$
from Section 7.2. We observe that, by Proposition 7.6, the application of the $1$st Eilenberg subcomplex functor $E_{1}$ is superfluous for this localization, for we are no longer working within the confines of $1$-reduced simplicial sets. In other words, left Bousfield localizing ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
$B_{M^{-1}\mathbb{Z}}=\left\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}$
yields the same model category, so we equivalently localize ${\mathcal{SS}\mathrm{ets}}$ at the set of maps $B_{M^{-1}\mathbb{Z}}$ instead.
10.2. The family of model structures on simplicial sets
For a multiplicative subset $M$ of $\mathbb{Z}$, we describe the left Bousfield localization of the model category ${\mathcal{SS}\mathrm{ets}}$ at the set of maps
$B_{M^{-1}\mathbb{Z}}=\left\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\right\}.$
In particular, we characterize the $B_{M^{-1}\mathbb{Z}}$-local objects in ${\mathcal{SS}\mathrm{ets}}$ in terms of the following definition.
Definition 10.1.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A simplicial set $K$ is $M$-local if $K$ is a Kan complex and $|K|$ is an $M$-local space. When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational simplicial set.
Example 10.2.
By Proposition 7.6, the target of every map in $B_{M^{-1}\mathbb{Z}}$ is an $M$-local simplicial set. Note that Example 3.2 says that the target of every map in $B_{M^{-1}\mathbb{Z}}$ is a Kan complex.
The following proposition helps us transfer the rational homotopy theory from Section 9 to simplicial sets.
Proposition 10.3 ([Hir03, Proposition 1.1.11]).
If $A$ and $X$ are simplicial sets and $X$ is a Kan complex, then there is a weak homotopy equivalence
$\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(A,X\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|A|,|X|\right)$
that is natural in $A$ and in $X$.
By a series of characterizations, we show that a simplicial set is $B_{M^{-1}\mathbb{Z}}$-local if and only if it is $M$-local.
Proposition 10.4.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ be a simplicial set. The following statements are equivalent.
@@ 10.2. The family of model structures on simplicial sets
    The simplicial set $X$ is $B_{M^{-1}\mathbb{Z}}$-local.
  2. (2)
    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}),X\right)$
    is a weak homotopy equivalence.
  3. (3)
    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right)$
    is a weak homotopy equivalence.
  4. (4)
    The simplicial set $X$ is a Kan complex such that, for every $n\geq 2$, the map
    $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|X|)$
    is a weak homotopy equivalence.
  5. (5)
    The simplicial set $X$ is a Kan complex such that $|X|$ is an $M$-local space.
  6. (6)
    The simplicial set $X$ is $M$-local.
Hence, $X$ is $B_{M^{-1}\mathbb{Z}}$-local if and only if $X$ is $M$-local.
Proof.
Firstly, (1) and (2) are equivalent by the definition of a $B_{M^{-1}\mathbb{Z}}$-local object in ${\mathcal{SS}\mathrm{ets}}$. Then, (2) and (3) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}\left(\operatorname{Sing}(S^{n}),X\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 10.3. Now, (3) and (4) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})|,|X|\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},|X|)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(|\operatorname{Sing}(S^{n})|,|X|\right),}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.6. Also, (4) and (5) are equivalent by Proposition 9.7. Lastly, (5) and (6) are equivalent by the definition of an $M$-local simplicial set. ∎
Proposition 10.4 helps us describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}$ at the set of maps $B_{M^{-1}\mathbb{Z}}$.
Theorem 10.5.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets in which a map $f\colon X\rightarrow Y$ is
  1. (1)
    a weak equivalence if and only if, for every $M$-local simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,K)$
    is a weak homotopy equivalence; and
@@
A simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is $M$-local. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ is combinatorial, left proper, and simplicial with the structure from Example 5.2.
Theorem 10.5 partially extends Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets. A distinguished model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
Theorem 10.6.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ on the category ${\mathcal{SS}\mathrm{ets}}$ of simplicial sets in which a map $f\colon X\rightarrow Y$ is
@@
    a weak equivalence if and only if, for every rational simplicial set $K$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,K)$
    is a weak homotopy equivalence; and
@@
    a cofibration if and only if $f$ is a monomorphism.
A simplicial set $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is rational. The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ is combinatorial, left proper, and simplicial with the structure from Example 5.2.
11. The family of model structures on topological spaces
Quillen’s family of model structures on the category of $1$-reduced simplicial sets finds no counterpart for the category of $1$-connected spaces, for the latter category cannot be endowed with any model structure because it does not have all finite limits and colimits; this failure necessitates the use of $1$-reduced simplicial sets in Quillen’s model-categorical approach of rational homotopy theory.
To see that the category of $1$-connected spaces does not have all finite limits, consider the pullback square
[…]
where $f\colon\mathbb{R}^{2}\rightarrow\mathbb{R}$ is given by the formula $f(x,y)=x^{2}+y^{2}$, so $f^{-1}(1)=S^{1}$ is the unit circle in the plane. In this pullback square, we start with the contractible spaces $|\Delta[0]|$, $\mathbb{R}^{2}$, and $\mathbb{R}$, and we obtain $S^{1}$, which is not a $1$-connected space because $\pi_{1}(S^{1})\cong\mathbb{Z}$. Thus, we see that connectedness conditions on spaces do not interact well with limits and colimits, in contrast to skeletal conditions on simplicial sets.
However, now that we have a family of model structures on the category of simplicial sets, we continue our work by presenting its analog for the category $\mathcal{T}\mathrm{op}$ of topological spaces. This analog is produced by left Bousfield localizing the model structure on $\mathcal{T}\mathrm{op}$, which we recall in this section. Moreover, because we first characterized local topological spaces and then transferred this characterization to simplicial sets in this section, there is no additional work that we need to do in order to obtain a family of model structures on topological spaces.
@@ 11.1. The model category of topological spaces
We recall the model structure on $\mathcal{T}\mathrm{op}$ from [DS95, §8]. To describe it, we first need to introduce the following notions.
@@
    Given a sequence of subspace inclusions
    $X_{0}\rightarrow X_{1}\rightarrow\cdots$
    such that, for every $n\geq 0$, the pair $(X_{n+1},X_{n})$ is a relative CW complex, we say that the map
    $X_{0}\rightarrow\operatorname{colim}(X_{n})$
    is a generalized relative CW inclusion.
@@
    A Serre fibration is a continuous map $p\colon X\rightarrow Y$ that has the right lifting property with respect to the class of inclusions
    $\left\{A\times\{0\}\rightarrow A\times[0,1]\mid A\textrm{ is a CW complex}\right\}.$
    In other words, a Serre fibration is a continuous map that has the homotopy lifting property with respect to every CW complex.
Remark 11.2.
A continuous map is often said to be a fibration if it has the homotopy lifting property with respect to every space. This notion is stronger than that of a Serre fibration, and it is sometimes called a Hurewicz fibration to emphasize this distinction. Every Hurewicz fibration is a Serre fibration; the converse implication finds a pathological counterexample in [Bro66, §1], but it holds for maps of CW complexes [Cau92].
Example 11.3.
Every space $X$ is Serre fibrant: the unique map from $X$ to the point is a Serre fibration.
We are now ready to describe the model category of topological spaces.
Theorem 11.4 ([DS95, Propositions 8.3 and 8.9]).
@@
The model category of topological spaces is simplicial with the following structure.
Proposition 11.5 ([Hir03, Example 9.1.15]).
@@
    the function complex $\operatorname{Map}_{{\mathcal{T}\mathrm{op}}}\left(X,Y\right)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{T}\mathrm{op}}}\left(X\times|\Delta[n]|,Y\right)$;
@@
The model category of topological spaces is also cofibrantly generated; its generating cofibrations arise from geometrically realizing those for the model category of simplicial sets.
Proposition 11.6 ([Hir03, Example 11.1.8]).
The model category ${\mathcal{T}\mathrm{op}}$ is cofibrantly generated.
@@
    A generating set of cofibrations for ${\mathcal{T}\mathrm{op}}$ is
    $|I|=\left\{|i_{n}|\colon|\partial\Delta[n]|\rightarrow|\Delta[n]|\mid n\geq 0\right\}.$
@@
    A generating set of acyclic cofibrations for ${\mathcal{T}\mathrm{op}}$ is
    $|J|=\left\{|j_{n,k}|\colon|\Lambda[n,k]|\rightarrow|\Delta[n]|\mid n\geq 1\textrm{ and }0\leq k\leq n\right\}.$
The model category of topological spaces differs from that of simplicial sets in two noteworthy respects.
@@
    It is left proper [DS95, Theorem 13.1.10], but not by Proposition 3.12, for not every space is cofibrant.

### v2

@@ 9.2. The construction of local spheres
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. To salvage our function complex characterization of $M$-local spaces in the absence of a universal property for $M$-localizatrion, we need the details of the CW complex construction of the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ from [22, Chapter 9(a)].
We enumerate $M\cap\mathbb{N}=\{m_{1},m_{2},m_{3},\dots\}.$ We denote the $(n+1)$-disk by $D^{n+1}$. Then, $S^{n}_{M^{-1}\mathbb{Z}}$ is the pushout
$\coprod\limits_{j=1}^{\infty}(S^{n})_{j}$$\bigvee\limits_{i=0}^{\infty}(S^{n})_{i}$$\coprod\limits_{j=1}^{\infty}(D^{n+1})_{j}$$S^{n}_{M^{-1}\mathbb{Z}},$$\scriptstyle{h}$
$S^{n}\xrightarrow{\textrm{pinch}}S^{n}\vee S^{n}\xrightarrow{\operatorname{id}_{S^{n}}\vee(-m_{j})}(S^{n})_{j-1}\vee(S^{n})_{j}\xrightarrow{\textrm{inclusion}}\bigvee\limits_{i=0}^{\infty}(S^{n})_{i}.$
The map $(-m_{j})$ is a self-map of $S^{n}$ that has degree $(-m_{j})$. In other words, the $1$-connected CW complex
@@
is the mapping telescope of the sequence of maps $S^{n}\xrightarrow{m_{1}}S^{n}\xrightarrow{m_{2}}S^{n}\xrightarrow{m_{3}}\cdots$.
9.3. A characterization of rational spaces
We use the construction of $S^{n}_{M^{-1}\mathbb{Z}}$ to prove a key result.
Proposition 9.6.
@@ 9.3. A characterization of rational spaces
Given a map $f\colon S^{n}\rightarrow Y$, we know that $[f]$ is an element of the $M^{-1}\mathbb{Z}$-module $\pi_{n}(Y)$. Thus, there exists a unique element $[u]$ of $\pi_{n}(Y)$ such that $f]=m_{1}\cdot[u]$, and the representative $u\colon S^{n}\rightarrow Y$ of $[u]$ is unique up to homotopy. Using the notation in our construction of $S^{n}_{M^{-1}\mathbb{Z}}$, we extend $f$ from $S^{n}=(S^{n})_{0}$ to $(S^{n})_{0}\vee(S^{n})_{1}$ uniquely up to homotopy by setting $g|_{(S^{n})_{1}}=u$. As $f]-(m_{1}\cdot[u])=0$, we can further extend $f$ to
@@
Proposition 9.6 salvages our characterization of $M$-local spaces in terms of function complexes.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence.
Proof.
If $Y$ is $M$-local, then Proposition 9.6 implies that the restriction map is a weak equivalence. The converse implication is established as in the proof of Proposition 6.8. ∎
10. The family of model structures on spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We left Bousfield localize the model category $\mathcal{SS}\mathrm{ets}$} of simplicial sets to obtain a model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ in which the weak equivalences are the $M$-local homotopy equivalences in our extended sense and the fibrant objects are the $M$-local Kan complexes. In particular, we produce a model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ whose weak equivalences are the rational homotopy equivalences and whose fibrant objects are the rational Kan complexes. Thus, we generalize Quillen’s family of model structures to all spaces. Lastly, we show that our localization of spaces at rational homotopy does not coincide with Bousfield’s localization at rational homology as rational homotopy and rational homology do not agree beyond the $1$-connected setting.
10.1. The family of model structures on simplicial sets
Proposition 9.6 allows us to describe the left Bousfield localization of $\mathcal{SS}\mathrm{ets}$} at the set of maps
$T_{M^{-1}\mathbb{Z}}=\{\operatorname{Sing}(S^{n})\rightarrow\operatorname{Sing}(S^{n}_{M^{-1}\mathbb{Z}})\mid n\geq 2\}.$
To do so, we first formulate the analog of Definition 9.1 for simplicial sets.
Definition 10.1.
Let $M$ be a multiplicative subset of $\mathbb{Z}$.
  1. (1)
    A simplicial map $f$ is an $M$-local homotopy equivalence if $|f|$ is an $M$-local homotopy equivalence.
  2. (2)
    A simplicial set $K$ is $M$-local if $|K|$ is an $M$-local space.
  3. (3)
    An $M$-localization of $X$ is an $M$-local homotopy equivalence with $M$-local target $X\rightarrow X_{M^{-1}\mathbb{Z}}$.
Having salvaged our characterization of $M$-local spaces in terms of function complexes, we can replicate our proof of Proposition 7.9 to show that a simplicial set is $T_{M^{-1}\mathbb{Z}}$-local if and only if it is an $M$-local Kan complex. It remains to show that a $T_{M^{-1}\mathbb{Z}}$-local equivalence is precisely an $M$-local homotopy equivalence. To that end, we invoke a general result on fiberwise localizations for the fiberwise $M$-localization from Proposition 9.5.
Lemma 10.2.
[26, Theorem 6.1.3(3)] Let $M$ be a multiplicative subset of $\mathbb{Z}$. Every fiberwise $M$-localization map is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
Lemma 10.2 allows us to characterize the $T_{M^{-1}\mathbb{Z}}$-local equivalences in our desired fashion.
Proposition 10.3.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a simplicial map. The following statements are equivalent.
  1. (1)
    The map $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
  2. (2)
    The induced map on fiberwise $M$-localizations $f_{M^{-1}\mathbb{Z}}$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence.
  3. (3)
    The induced map on fiberwise $M$-localizations $f_{M^{-1}\mathbb{Z}}$ is a weak equivalence.
  4. (4)
    The induced map on fiberwise $M$-localizations $f_{M^{-1}\mathbb{Z}}$ is an $M$-local homotopy equivalence.
  5. (5)
    The map $f$ is an $M$-local homotopy equivalence.
Proof.
Fiberwise $M$-localization induces the commutative square
@@ 10.1. The family of model structures on simplicial sets
Firstly, (1) and (2) are equivalent because the vertical maps are $T_{M^{-1}\mathbb{Z}}$-local equivalences by Lemma 10.2. Then, (2) and (3) are equivalent by Remark 7.2. Now, (3) and (4) are equivalent by Example 9.2. Lastly, (4) and (5) are equivalent because the vertical maps are $M$-local homotopy equivalences by Proposition 9.5.
We arrive at our promised generalization of Quillen’s family of model structures to all spaces.
Theorem 10.4.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ on the category of simplicial sets in which a map $f$ is
@@
A simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ if and only if $X$ is an $M$-local Kan complex. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}$ is combinatorial, left proper, and simplicial with the structure from Example 5.2.
Again, a distinguished model structure in this family is that corresponding to the rational case $M^{-1}\mathbb{Z}=\mathbb{Q}$.
Corollary 10.5.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}$ on the category of simplicial sets in which
@@
    a map is a weak equivalence if and only if it is a rational homotopy equivalence; and
@@
    a simplicial set is fibrant if and only if it is a rational Kan complex.
Thus, our modern approach to Quillen’s model structures on $1$-reduced simplicial sets generalizes to all spaces and yields a model category that encodes the rational homotopy theory of non-simply connected spaces.
@@ 10.2. The model category of topological spaces
Quillen’s family of model structures on the category of $1$-reduced simplicial sets finds no counterpart for the category of $1$-connected spaces, for the latter category does not have all finite limits and colimits. For instance, consider the pullback
$S^{1}$$\mathbb{R}^{2}$$|\Delta[0]|$$\mathbb{R}$$\scriptstyle{f}$$\scriptstyle{1}$
where $f$ is given by the formula $f(x,y)=x^{2}+y^{2}$, so $f^{-1}(1)=S^{1}$ is the unit circle in the plane. In this pullback, we start with the contractible spaces $|\Delta[0]|$, $\mathbb{R}^{2}$, and $\mathbb{R}$ and get $S^{1}$, which has fundamental group $\pi_{1}(S^{1})\cong\mathbb{Z}$. However, now that we have a family of model structures on simplicial sets, we can produce its analog for the category $\mathcal{T}\mathrm{op}$ of topological spaces by left Bousfield localizing the model structure on $\mathcal{T}\mathrm{op}$ at our usual set.
We first recall the model structure on $\mathcal{T}\mathrm{op}$ from [18, §8]; the description is in terms of the following definition.
@@
    Given a sequence of inclusions $X_{0}\rightarrow X_{1}\rightarrow\cdots$ such that the pairs $(X_{n+1},X_{n})$ are relative CW complexes, we say that $X_{0}\rightarrow\operatorname{colim}(X_{n})$ is a generalized relative CW inclusion.
@@
    A Serre fibration is a map that has the right lifting property with respect to the class of inclusions $\{A\times\{0\}\rightarrow A\times[0,1]\mid A\textrm{ is a CW complex}\}$.
In other words, a Serre fibration is a map with the homotopy lifting property against all CW complexes.
Theorem 10.7 ([18, Propositions 8.3 and 8.9]).
@@
The model category of spaces is also cofibrantly generated and simplicial with the following structure.
Proposition 10.8 ([26, Example 11.1.8]).
The model category $\mathcal{T}\mathrm{op}$} is cofibrantly generated.
  1. (1)
    A generating set of cofibrations is $|I|=\{|i_{n}|\colon|\partial\Delta[n]|\rightarrow|\Delta[n]|\mid n\geq 0\}$.
  2. (2)
    A generating set of acyclic cofibrations is $|J|=\{|j_{n,k}|\colon|\Lambda[n,k]|\rightarrow|\Delta[n]|\mid n\geq 1\textrm{ and }0\leq k\leq n\}$.
Proposition 10.9 ([26, Example 9.1.15]).
@@
    the function complex $\operatorname{Map}_{{\mathcal{T}\mathrm{op}}}(X,Y)$ to be the simplicial set with $n$-simplices $\operatorname{Hom}_{{\mathcal{T}\mathrm{op}}}(X\times|\Delta[n]|,Y)$;
@@
The model category of topological spaces differs from that of simplicial sets in three noteworthy respects.
@@
    It is not combinatorial because the category $\mathcal{T}\mathrm{op}$} is not locally presentable [1, Example 1.18(5)].
@@
    It is left proper [18, Theorem 13.1.10], but not by Proposition 3.12, for not every space is cofibrant.
@@
    Every topological space is Serre fibrant, whereas not every simplicial set is a Kan complex.
Nevertheless, the geometric realization-singular complex adjunction is a Quillen equivalence [32, Chapter I.4, Example 2], which renders simplicial sets a model for spaces. Still, the first difference raises the question of whether $\mathcal{T}\mathrm{op}$} can be localized. We note that $\mathcal{T}\mathrm{op}$} shares with $\mathcal{SS}\mathrm{ets}$} the feature of being a cellular model category [26, Proposition 12.1.4]. This property is a different way of strengthening cofibrant generation, and it lets us localize $\mathcal{T}\mathrm{op}$}, for Theorem 7.3 holds if we replace “combinatorial” with “cellular” [26, Theorem 4.1.1].
10.3. The family of model structures on topological spaces
Corollary 9.7 and Proposition 10.3 describe the left Bousfield localization of $\mathcal{T}\mathrm{op}$ at the set of maps $\{S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}\mid n\geq 2\}$ with no additional work needed.
Theorem 10.10.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{T}\mathrm{op}}$ on the category of topological spaces in which a map $f$ is